OpenStudy (smartyprincess):

P=? 3p/6p+7=39 I think it's 8. Anyone agree?

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \frac{3p}{6p+7}=39 }\) or \(\color{black}{\displaystyle \frac{3p}{6p}+7=39 }\)

OpenStudy (smartyprincess):

first one

OpenStudy (smartyprincess):

I know the process on how to figure it out just need an aproval :)

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \frac{3p}{6p+7}=39 }\) \(\color{black}{\displaystyle \frac{3p}{6p+7}(6p+7)=39(6p+7) }\)

OpenStudy (solomonzelman):

multiply on both sides times the quantity in the denominator, so that you can get rid of the fraction, first.

OpenStudy (solomonzelman):

Can you simplify and expand on both sides?

OpenStudy (smartyprincess):

well it's not exactly a fraction it's dividing

OpenStudy (smartyprincess):

we will have to subtract 7 from 39 and 6p+7 and get 6p=32

OpenStudy (smartyprincess):

6 divided into 6p is p and 32 ÷ 6 = 5.3

OpenStudy (mathmale):

Once you have found a value of p, please check your answer yourself by substituting it into the original equation. Is that equation now true? If so, your "solution" is correct. Be really careful with those parentheses; they are not optional (unless you want confused helpers).

OpenStudy (smartyprincess):

if we replace 8 into the equation we get : 3•8=24 ________ 6 •8=48+7=55 24/55

OpenStudy (smartyprincess):

55 ÷ 24

OpenStudy (mathmale):

You think p=8? Please subst. 8 for p in the original equation. Is that equation now true?

OpenStudy (smartyprincess):

not sure

OpenStudy (smartyprincess):

I'm very confused

OpenStudy (mathmale):

Not sure about what? Confused about what? Please explain. Originally you wanted to know whether your solution, p=8, is correct; I am explaining how you can check it easily by yourself.

OpenStudy (smartyprincess):

oh I see

OpenStudy (mathmale):

\[\color{black}{\displaystyle \frac{3p}{6p+7}=39 }\]

OpenStudy (smartyprincess):

\[3p/6p+7=39 \]

OpenStudy (smartyprincess):

39-7=32

OpenStudy (smartyprincess):

i have to go now sorry

OpenStudy (mathmale):

Wait, please. Are you SURE you want 3p/6p+7 = 32? I don't agree, and neither does Solomon.

OpenStudy (mathmale):

Come back when you have adequate time to finish this.

OpenStudy (smartyprincess):

I will schedule a get together you could continue putting explanations in here

OpenStudy (smartyprincess):

ok thats fine

OpenStudy (smartyprincess):

get togthere with my teacher

OpenStudy (mathmale):

No. As a matter of principle I want to work with you while you're here and participating.

OpenStudy (smartyprincess):

ok

jimthompson5910 (jim_thompson5910):

Here's a similar example. Let's say we had the equation \(\Large x+2 = 6\) and I claim that \(\Large x = 8\) is the solution How can I test this claim? By following what @mathmale is suggesting. Replace every copy of x with 8 and see if you get a true equation. Let's do that \[\Large x+2 = 6\] \[\Large {\color{red}{x}}+2 = 6\] \[\Large {\color{red}{8}}+2 = 6\] \[\Large 10 = 6\] The last equation is false, so `x+2=6` is false when x = 8. Therefore, x = 8 is NOT the solution here.